OpenStudy (anonymous):

A sledgehammer weighs 18 pounds on Earth and 4 pounds on the planet Namar. Which of the following proportions could you use to determine a 100–pound earthling's weight on Namar? A. 18/4 = 100/w B. 4/18 = w/100 C. 18/4 = w/100 D. Both A & B are correct.

OpenStudy (anonymous):

Lost & confused.. >.< Please help.

OpenStudy (anonymous):

@NPTWEEZY @marissalovescats @loujoelou @kelliegirl33 @heather040200 @goformit100

OpenStudy (anonymous):

18/4 = 100/x

OpenStudy (anonymous):

wait.....A and B are correct

OpenStudy (anonymous):

Yeah, thats what I thought at first if I should choose D. & Can you please explain? @kelliegirl33

OpenStudy (anonymous):

you can work them out and see.... 18/4 = 100/w (cross multiply) 18w = 400 w = 22.22 4/18 = w/100 (cross multiply) 18w = 400 w = 22.22 what you have to make sure of, is that whatever you put on the numerator of one, you put on the numerator of the other....same goes for the denominators... lbs on earth/lbs on namar....earth lbs on top (numerator), lbs on namar on bottom (denominator). Now if you put lbs on namar on top, you have to put lbs of namar on the numerator of the other one also. They have to line up. Does that make sense ?

OpenStudy (anonymous):

Yes I understand, can you please help me with the question I just tagged you in? :)